OpenStudy (anonymous):

2. (Growth & Decay) A certain radioactive material has a half-life of 28 hours. A scientist finds 38.6 grams of this material. (a) The scientist plans to do an experiment with this material, but it takes 7 hours to set up the equipments for the lab. By the time that the laboratory is ready, how much of the material will there be left? (b) If the scientist needs at least 34 grams of the material to perform a successful experiment, how much time does he have to set up the equipments? (c) If the scientist could have found this material 3 days (72 hours) ago, what would be the mass

OpenStudy (anonymous):

of the material at that time?

OpenStudy (anonymous):

first one, answer is \[38.6\left(\frac{1}{2}\right)^{\frac{7}{28}}\] \[38.6\left ( \frac{1}{2}\right)^{\frac{1}{4}}\] now a calculator